### v4

@@ 2.1 Data, Definitions, and Descriptive Statistics
Research activity–defined by publications–varies considerable by school and over time. Column 3 in Table 1 shows that roughly 37-56% of 1994-2017 graduates from economics departments published at least one DI paper in the six years following degree receipt. As can be seen in column 6, research activity rates are lower for graduates of economics-related programs.
@@ 3 Research Output
Many graduates fail to produce a single publication, a finding reported earlier in Conley and Önder (2014). Figure 4 shows something similar for our graduates. Among graduates of economics and related programs, fewer than 60% ever place a paper in a DI journal; only around 45% place two. These long-run success rates are only a little higher among economics department graduates. T6 publications are rare: only about a quarter of graduates place a paper in a top-6 outlet and only around 15% manage two in T6. The likelihood of a T6 publication flattens a little sooner than does the odds of placing a paper in DI journals. T6 success goes higher for economics graduates, though still fails to clear 30%.
@@ 3.1 Gender Gaps
Research activity profiles by time since degree, plotted by gender in Figure 8 show male and female graduates to be similarly active in the first few years post-PhD. By year five, however, female DI activity rates have crested while male DI activity rates continue to climb. This leads to a gap in activity rates that begins to close only 15 years after graduation. Consistent with much lower T6 activity, gender gaps in T6 publication rates are smaller than those for DI, and appear to close sooner than those for DI publications. As a proportion of corresponding male activity rates, however, the largest gaps (observed between post-PhD years 5-10) are similar for DI and T6 publications. [^8] Proportional gaps are $0.22$ for DI and $0.21$ for T6.
@@ 4.1 Advisor Attributes and Student Success
In models entering advisor attributes one at a time, graduates advised by advisors who publish more, advisors with more advisees, and advisors with more successful former students have greater publication success. The first column of Panel A in Table 2 shows, for instance, that graduates see $0.12$ more DI publications, on average, when advising team research increases by one. This is a gain of roughly 7% (the corresponding Poisson coefficient appears in column 5). As can be seen in column 3, advisor research success appears to boost advisees’ T6 publications much less, though effects on advisees’ T6 publications in percentage terms (reported in column 7) are larger and approach 9%. The estimates in the second row of Panel A show that graduates advised by super researchers generate more publications. Super-research effects are $0.69$ and $0.21$ for DI and T6 levels, respectively, and $0.48$ and $0.68$ in percent. [^14] García-Suaza, Otero and Winkelmann (2020); Hilmer and Hilmer (2009) likewise find a strong association between advisor research productivity and economics advisees’ post-PhD publishing success.
The estimates in Panel B of Table 2 suggest that an advising team’s advising load is (mostly) a weaker predictor of student research success than the team’s own research record. Patterns here are similar to those in Panel A, however, with consistently positive effects: prolific advisors (in terms of numbers of past students) have advisees that see greater post-graduation research success. As with the estimated advisor-research effects in Panel A, effects on student output in levels are larger for DI than T6, while effects in percentage terms are larger than the corresponding levels coefficients for T6. Super advising-loaded advisors have substantially more successful students than do advisors with fewer advisees.
Among the three continuous advisor attributes examined in Table 2, the estimates in Panel C of the table show that average past student success is the best predictor of current student success. Increasing an advisor team’s average past student DI publications by 1, for instance, is associated with $0.43$ more current student DI publications, a 26% increase. Increasing the average of past student DI publications is predicted to yield a 36% increase in current student T6 publications. Again, coefficients on super dummies indicating the strongest 10% of advisors based on their advisees’ past success are even larger.
@@
Multivariate models generate dampened effects of advising load on student research success. Interestingly, however, the estimates in Tables 2 and Tables show no evidence of a productivity-diminishing advising burden: advisees guided by advisors carrying a heavy load do not appear to suffer when their advisors’ attention is divided more finely. Figure 13 highlights the weakly positive relationship between advising load and student research success. This figure plots average student success for all 1989-2023 graduates against the number of students advised, tabulated for the sample of advisors who advised at least 30 students between 1989-2023.151515Let $S_{j}$ denote the set of advisor $j$’s students, indexed by $i$, who graduated 1989-2023. Overall student success in Figure 13, denoted $PS_{j}$ for advisor $j$, is defined as: $PS_{j}\equiv\frac{1}{|S_{j}|}\sum_{\{i\in S_{j}\}}R_{i},$ (6) where $R_{i}$ counts DI publications by student $i$ who graduated in cohort $c$ in years $c+1$ to $c+6$. The x-axis is a log scale, so the slope of the fitted line implies that a 10% increase in advising load is associated with $0.036$ additional post-PhD publications on average, with considerable variation around this.
The figure is also consistent with the robustly positive duper advising load coefficient estimates reported in Table Tables in suggesting an outsize role for exceptionally prolific advisors (measured by their advising loads). The five most prolific advisors (Acemoglu, Cutler, Card, Katz, and Shleifer) are labeled in the figure; these lie to the Northeast of the mass of points marking other advisors. Figure 13 is complemented by Appendix Table LABEL:t:league, which lists the 180 advisors whose students have the strongest DI publication records, ranked by student success. The table of top advisors, which includes number of advisees and advisor rank by advising load, shows many heavily-loaded advisors with strong students. [^16] The table is truncated from below at prolific advisor and Nobel laureate Robert E. Lucas, whose many advisees averaged just over 1 publication post-PhD. The figure plots data for all 235 advisors in our data who advised at least 30 students. Of course, the correlations and regression estimates discussed in this section may reflect student selection or sorting as well as causal effects. We return to this point following an examination of student-driven aspects of the advising relationship.
@@ 4.2 Advisor-Advisee Coauthoring and Research Affinity
Unsurprisingly, coauthoring of any kind is far more common after degree receipt than before (since PhD students must publish to coauthor). For all cohorts, coauthoring with advisors is more common than coauthoring with classmates. But both sorts of coauthoring are on the rise. Among the most recent cohorts for which data appear in Figure 14, advisor-advisee coauthoring rates hit 20%, while classmate coauthoring reaches roughly 15%. [^18] Sarsons et al. (2021) examines gender differences in credit attribution for joint work, finding statistically insignificant gender gaps in post-tenure publications and citations to jointly-authored research.
@@
Our measure of thesis citations to advisors is imperfect and relies in large part on algorithms that read thesis PDFs. [^19] Thesis citations are not indexed by the Web of Science or EconLit. We therefore search thesis PDFs for all article titles of advisors’ publication lists as detailed in the data appendix. Advisor affiliation is also approximate in our data, relying on author affiliations attached to publications in EconLit. Still, the view that the resulting citation rates are informative is supported by Figure 16. This figure tracks the number of students advised and student citations to advisors for advisors that change affiliations. For 367 fixed-super advisors (defined as an advisor whose advising load falls in the upper decile of the advisee distribution for any cohort), the figure counts advisees at the former affiliation and new affiliation separately. A transitioning advisor’s advising load at their previous affiliation falls sharply at the time of a move, plateauing close to zero four years later. At the same time, the number of advisees advised by transitioning advisors at the new institution increases rapidly. A similar pattern appears in the number of graduates that cite transitioning advisors’ work: their number decreases at a transitioning advisor’s previous institution while it increases at the new one (considering all students at origin and destination schools, not just graduates advised by transitioning advisors). [^20] Our work here is in the spirit of Lerner et al. (2024), which traces patent citations to transitioning researchers’ papers as they move across institutions. These patterns suggest advisees citations capture a decline in intellectual influence at an advisor’s old institution and rising influence at the new one.
@@
When included in a regression model like equation (5), variables that indicate coauthoring with advisors and classmates before or in graduation years are unrelated to PhD students’ immediate post-graduation publication success. Estimated advisor-coauthoring effects on DI publications, reported in the first four columns of Table A2, are small and not significantly different from zero. Classmate-coauthoring coefficient estimates, reported in columns 2 and 4 of the table, are likewise not significantly different from zero, though estimated much less precisely than the corresponding advisor-coauthoring effects.
@@
Graduates whose thesis cites an advisor, by contrast, are estimated to see $0.22$ more DI publications and $0.07$ more T6 publications, estimated gains that are significantly different from zero. The Poisson estimates in columns 6 and 8 of Table A2 show that these gains amount to publication increases of 16% and 27%, respectively. The strongest predictor of post-PhD publication success is a dummy variable indicating graduates with any pre-PhD publication; this is a control necessitated by the fact that our definition of coauthoring requires publication and precocious pre-degree publishers are more likely to publish after leaving the nest.
@@
As can be seen in the first three rows of Table A2, estimated advisor effects generated by the extended version of equation (5) with coauthoring and affinity variables are similar to those in the first three rows of Table 3, with significant estimates of coefficients on advisor research and past student success. Replacing continuous advisor attributes with dummies indicating fixed-duper advisors generates the estimates in Table LABEL:t:_Coauthoring_Fixed_Duper. In this model, fixed-duper dummies indicate advisors with a given attribute (e.g., number of advisees) ever coded as duper. This specification is motivated by a view of exceptional advisor performance as a time-invariant attribute. The resulting estimates are broadly consistent with those in Tables Tables and A2, suggesting that the most important advisor attributes are advisor research and past student success. In contrast with the estimates in Table A2, however, estimated fixed-duper advising load coefficients are larger–around $0.1$–and significantly different from zero for graduates’ T6 publications. It’s noteworthy, however, that regardless of specification, student success is hard to predict: for all models reported in Tables Tables-LABEL:t:_Coauthoring_Fixed_Duper, $R^{2}$s are no bigger than .14 for DI publications and below .1 for T6.
@@ 5.1 Aggregation IV
Importantly, two-stage least squares (2SLS) estimates based on restriction (8) are not confounded by within-school sorting of students to advisors. Suppose, for instance, that within departments, the best students seek out advisors with strong research records, but advisor research success generates no payoff in terms of advisee research success. In this scenario, estimates like those in Table A2 are likely to be positive: the fact that productive students seek out productive advisors engenders positive omitted variables bias. Yet, such within-school sorting leaves average student success by cohort and school unchanged, making the aggregate student research production function a better guide to causal advisor effects than equation (5) estimated using data on individual graduates. Because the aggregate model controls for cohort and school effects, restriction (8) identifies causal effects by exploiting cross-cohort changes in average advising features within schools. These changes are due both to advisor transitions between schools and evolving within-department changes in incumbent advisor features. Elite departments compete to attract top scholars and prolific advisors, hoping (among other hiring goals) for an immediate boost in advising horsepower. As a suggestive exploration of the role played by advisor transitions, Figure 17 plots average research output for graduates at schools losing and receiving transitioning fixed-duper advisors as determined by their past student success. This figure is constructed in a manner similar to that used to construct Figure 16. We opt here for fixed-duper rather than fixed-super advisor transitions because the former generate larger student-level estimates in Table LABEL:t:_Coauthoring_Fixed_Duper. In contrast with Figure 16, which shows advising load and citation changes consistent with transitioning advisors having an impact on affected schools, the pattern of student success traced in Figure 17 shows little evidence of gains in student research success at departments bolstered by the arrival of a fixed-duper advisor.
@@
Consistent with Figure 17, 2SLS estimates using 49 dummy instruments provide less evidence that advising features matter than do the corresponding OLS estimates. 2SLS estimates with all features instrumented, reported in the first four columns of Table LABEL:t:_2SLS, generates statistically significant estimates only for the impact of advisor research. Statistically significant 2SLS estimates of the impact of advisor research, on the order of $0.16$ for graduate DI publications and $0.047$ for graduate T6 publications, exceed the corresponding OLS estimates in columns 1-4 of Table A2 but are less precise. 2SLS estimates instrumenting only the three advisor attributes generate large, precisely-estimated pre-grad publication effects similar those in Table A2.
Estimates of coauthoring and advisor-citation effects in Table LABEL:t:_2SLS are so imprecise they should be seen as uninformative, though large advisor-coauthoring coefficients are marginally significant for graduates’ T6 publications in column 3. On balance, this table does little to bolster claims for a strong relationship between advisor attributes and student success. Most notably, the coefficient on advising load is a reasonably-precisely estimated statistical zero, while the coefficient on past student success is precise enough that some effects of the magnitude seen in Table A2 would be marginally significantly different from zero. The strongest evidence for causal effects emerges for advisors’ own research output, highlighting the critical role of advisor research success in shaping their advisees post-PhD research outcomes. Factors that seem to matter in Tables A2 and LABEL:t:_Coauthoring_Fixed_Duper fail to generate consistent or precisely-estimated changes in average PhD student research success.

### v5

@@ 2.1 Data, Definitions, and Descriptive Statistics
Research activity–defined by publications–varies considerable by school and over time. Column 3 in Table 1 shows that roughly 38-58% of 1994-2017 graduates from economics departments published at least one DI paper in the six years following degree receipt. As can be seen in column 6, research activity rates are lower for graduates of economics-related programs.
@@ 3 Research Output
Many graduates fail to produce a single publication, a finding reported earlier in Conley and Önder (2014). Figure 4 shows something similar for our graduates. Among graduates of economics and related programs, fewer than 60% ever place a paper in a DI journal; only around 45% place two. These long-run success rates are only a little higher among economics department graduates. T6 publications are rare: only 30% of graduates place a paper in a top-6 outlet and only around 20% manage two in T6. The likelihood of a T6 publication flattens a little sooner than does the odds of placing a paper in DI journals, again with slightly higher success rates for economics department graduates.
@@ 3.1 Gender Gaps
Research activity profiles by time since degree, plotted by gender in Figure 8 show male and female graduates to be similarly active in the first few years post-PhD. By year five, however, female DI activity rates have crested while male DI activity rates continue to climb. This leads to a gap in activity rates that begins to close only 15 years after graduation. Consistent with much lower T6 activity, gender gaps in T6 publication rates are smaller than those for DI, and appear to close sooner than those for DI publications. As a proportion of corresponding male activity rates, however, the largest gaps (observed between post-PhD years 5-10) are similar for DI and T6 publications. [^8] Proportional gaps are $0.23$ for DI and $0.20$ for T6.
@@ 4.1 Advisor Attributes and Student Success
In models entering advisor attributes one at a time, graduates advised by advisors who publish more, advisors with more advisees, and advisors with more successful former students have greater publication success. The first column of Panel A in Table 2 shows, for instance, that graduates see $0.12$ more DI publications, on average, when advising team research increases by one. This is a gain of roughly 7% (the corresponding Poisson coefficient appears in column 5). As can be seen in column 3, advisor research success appears to boost advisees’ T6 publications much less, though effects on advisees’ T6 publications in percentage terms (reported in column 7) are larger and approach 9%. The estimates in the second row of Panel A show that graduates advised by super researchers generate more publications. Super-research effects are $0.70$ and $0.24$ for DI and T6 levels, respectively, and $0.47$ and $0.67$ in percent. [^14] García-Suaza, Otero and Winkelmann (2020); Hilmer and Hilmer (2009) likewise find a strong association between advisor research productivity and economics advisees’ post-PhD publishing success.
The estimates in Panel B of Table 2 suggest that an advising team’s advising load is (mostly) a weaker predictor of student research success than the team’s own research record. Patterns here are similar to those in Panel A, however, with consistently positive effects: prolific advisors (in terms of numbers of past students) have advisees that see greater post-graduation research success. As with the estimated advisor-research effects in Panel A, effects on student output in levels are larger for DI than T6, while effects in percentage terms are larger for T6 than for DI. Super advising-loaded advisors have substantially more successful students than do advisors with fewer advisees.
Among the three continuous advisor attributes examined in Table 2, the estimates in Panel C of the table show that average past student success is the best predictor of current student success. Increasing an advisor team’s average past student DI publications by 1, for instance, is associated with $0.43$ more current student DI publications, a 25% increase. Increasing the average of past student DI publications is predicted to yield a 35% increase in current student T6 publications. Again, coefficients on super dummies indicating the strongest 10% of advisors based on their advisees’ past success are even larger.
@@
Multivariate models generate dampened effects of advising load on student research success. Interestingly, however, the estimates in Tables 2 and Tables show no evidence of a productivity-diminishing advising burden: advisees guided by advisors carrying a heavy load do not appear to suffer when their advisors’ attention is divided more finely. Figure 13 highlights the weakly positive relationship between advising load and student research success. This figure plots average student success for all 1989-2023 graduates against the number of students advised, tabulated for the sample of advisors who advised at least 30 students between 1989-2023.151515Let $S_{j}$ denote the set of advisor $j$’s students, indexed by $i$, who graduated 1989-2023. Overall student success in Figure 13, denoted $PS_{j}$ for advisor $j$, is defined as: $PS_{j}\equiv\frac{1}{|S_{j}|}\sum_{\{i\in S_{j}\}}R_{i},$ (6) where $R_{i}$ counts DI publications by student $i$ who graduated in cohort $c$ in years $c+1$ to $c+6$. The x-axis is a log scale, so the slope of the fitted line implies that a 10% increase in advising load is associated with $0.035$ additional post-PhD publications on average, with considerable variation around this.
The figure is also consistent with the robustly positive duper advising load coefficient estimates reported in Table Tables in suggesting an outsize role for exceptionally prolific advisors (measured by their advising loads). The five most prolific advisors (Acemoglu, Cutler, Card, Katz, and Shleifer) are labeled in the figure; these lie Northeast of the mass of points marking other advisors. Figure 13 is complemented by Appendix Table LABEL:t:league and Appendix Figure LABEL:f:rankfig. The table lists the 189 most heavily-loaded advisors, ranked by their students’ average DI publishing success. This table includes advisor rank by student T6 publications, number of advisees, and advisor rank by advising load. [^16] The table is truncated from below at prolific advisor and Nobel laureate Robert E. Lucas, whose many advisees averaged just over 1 publication post-PhD. The figure plots data for all 236 advisors who advised at least 30 students. Appendix Figure LABEL:f:rankfig plots advisor rank based on student T6 publications against advisor rank based on student DI publications. These measures are highly correlated, with a regression slope of $0.81$ and an $R^{2}$ of $0.66$. Of course, the correlations and regression estimates discussed in this section may reflect student selection or sorting as well as causal effects. We return to this point following an examination of student-driven aspects of the advising relationship.
@@ 4.2 Advisor-Advisee Coauthoring and Research Affinity
Unsurprisingly, coauthoring of any kind is far more common after degree receipt than before (since PhD students must publish to coauthor). For all cohorts, coauthoring with advisors is more common than coauthoring with classmates. But both sorts of coauthoring are on the rise. Among the most recent cohorts for which data appear in Figure 14, advisor-advisee coauthoring rates hit 20%, while classmate coauthoring reaches roughly 17%. [^18] Sarsons et al. (2021) examines gender differences in credit attribution for joint work, finding statistically insignificant gender gaps in post-tenure publications and citations to jointly-authored research.
@@
Our measure of thesis citations to advisors is imperfect and relies in large part on algorithms that read thesis PDFs. [^19] Thesis citations are not indexed by the Web of Science or EconLit. We therefore search thesis PDFs for all article titles of advisors’ publication lists as detailed in the data appendix. Advisor affiliation is also approximate in our data, relying on author affiliations attached to publications in EconLit. Still, the view that the resulting citation rates are informative is supported by Figure 16. This figure tracks the number of students advised and student citations to advisors for advisors that change affiliations. For 364 fixed-super advisors (defined as an advisor whose advising load falls in the upper decile of the advisee distribution for any cohort), the figure counts advisees at the former affiliation and new affiliation separately. A transitioning advisor’s advising load at their previous affiliation falls sharply at the time of a move, plateauing close to zero four years later. At the same time, the number of advisees advised by transitioning advisors at the new institution increases rapidly. A similar pattern appears in the number of graduates that cite transitioning advisors’ work: their number decreases at a transitioning advisor’s previous institution while it increases at the new one (considering all students at origin and destination schools, not just graduates advised by transitioning advisors). [^20] Our work here is in the spirit of Lerner et al. (2024), which traces patent citations to transitioning researchers’ papers as they move across institutions. These patterns suggest advisees citations capture a decline in intellectual influence at an advisor’s old institution and rising influence at the new one.
@@
When included in a regression model like equation (5), variables that indicate coauthoring with advisors and classmates before or in graduation years are unrelated to PhD students’ immediate post-graduation publication success. Estimated advisor-coauthoring effects on DI publications, reported in the first four columns of Table Tables, are small and not significantly different from zero. Classmate-coauthoring coefficient estimates, reported in columns 2 and 4 of the table, are on the margin or not significantly different from zero, though estimated much less precisely than the corresponding advisor-coauthoring effects.
@@
Graduates whose thesis cites an advisor, by contrast, are estimated to see $0.23$ more DI publications and $0.07$ more T6 publications, estimated gains that are significantly different from zero. The Poisson estimates in columns 6 and 8 of Table Tables show that these gains amount to publication increases of 15% and 22%, respectively. The strongest predictor of post-PhD publication success is a dummy variable indicating graduates with any pre-PhD publication; this is a control necessitated by the fact that our definition of coauthoring requires publication and precocious pre-degree publishers are more likely to publish after leaving the nest.
@@
As can be seen in the first three rows of Table Tables, estimated advisor effects generated by the extended version of equation (5) with coauthoring and affinity variables are similar to those in the first three rows of Table 3, with significant estimates of coefficients on advisor research and past student success. Replacing continuous advisor attributes with dummies indicating fixed-duper advisors generates the estimates in Table LABEL:t:_Coauthoring_Fixed_Duper. In this model, fixed-duper dummies indicate advisors with a given attribute (e.g., number of advisees) ever coded as duper. This specification is motivated by a view of exceptional advisor performance as a time-invariant attribute. The resulting estimates are broadly consistent with those in Tables Tables and Tables, suggesting that the most important advisor attributes are advisor research and past student success. In contrast with the estimates in Table Tables, however, estimated fixed-duper advising load coefficients are larger–around $0.1$–and significantly different from zero for graduates’ T6 publications. It’s noteworthy, however, that regardless of specification, student success is hard to predict: for all models reported in Tables Tables-LABEL:t:_Coauthoring_Fixed_Duper, $R^{2}$s are no bigger than .14 for DI publications and below .1 for T6.
@@ 5.1 Aggregation IV
Importantly, two-stage least squares (2SLS) estimates based on restriction (8) are not confounded by within-school sorting of students to advisors. Suppose, for instance, that within departments, the best students seek out advisors with strong research records, but advisor research success generates no payoff in terms of advisee research success. In this scenario, estimates like those in Table Tables are likely to be positive: the fact that productive students seek out productive advisors engenders positive omitted variables bias. Yet, such within-school sorting leaves average student success by cohort and school unchanged, making the aggregate student research production function a better guide to causal advisor effects than equation (5) estimated using data on individual graduates. Because the aggregate model controls for cohort and school effects, restriction (8) identifies causal effects by exploiting cross-cohort changes in average advising features within schools. These changes are due both to advisor transitions between schools and evolving within-department changes in incumbent advisor features. Elite departments compete to attract top scholars and prolific advisors, hoping (among other hiring goals) for an immediate boost in advising horsepower. As a suggestive exploration of the role played by advisor transitions, Figure 17 plots average research output for graduates at schools losing and receiving transitioning fixed-duper advisors as determined by their past student success. This figure is constructed in a manner similar to that used to construct Figure 16. We opt here for fixed-duper rather than fixed-super advisor transitions because the former generate larger student-level estimates in Table LABEL:t:_Coauthoring_Fixed_Duper. In contrast with Figure 16, which shows advising load and citation changes consistent with transitioning advisors having an impact on affected schools, the pattern of student success traced in Figure 17 shows little evidence of gains in student research success at departments bolstered by the arrival of a fixed-duper advisor.
@@
Consistent with Figure 17, 2SLS estimates using 49 dummy instruments provide less evidence that advising features matter than do the corresponding OLS estimates. 2SLS estimates with all features instrumented, reported in the first four columns of Table LABEL:t:_2SLS, generates statistically significant estimates only for the impact of advisor research. Statistically significant 2SLS estimates of the impact of advisor research, on the order of $0.16$ for graduate DI publications and $0.047$ for graduate T6 publications, exceed the corresponding OLS estimates in columns 1-4 of Table Tables but are less precise. 2SLS estimates instrumenting only the three advisor attributes generate large, precisely-estimated pre-grad publication effects similar those in Table Tables.
Estimates of coauthoring and advisor-citation effects in Table LABEL:t:_2SLS are so imprecise they should be seen as uninformative. On balance, this table does little to bolster claims for a strong relationship between advisor attributes and student success. Most notably, the coefficient on advising load is a reasonably-precisely estimated statistical zero and the coefficient on past student success is smaller than the effects seen in Table Tables and imprecisely estimated. The strongest evidence for causal effects emerges for advisors’ own research output, highlighting the critical role of advisor research success in shaping their advisees post-PhD research outcomes. Factors that seem to matter in Tables Tables and LABEL:t:_Coauthoring_Fixed_Duper fail to generate consistent or precisely-estimated changes in average PhD student research success.